### 12 Dec 2023 v1

@@ Abstract
The neutral kaons are eigenstates of strong interaction, oscillating ($K^{0}\leftrightarrow\bar{K}^{0}$) due to weak interaction. The weak eigenstates are the long-lived ‘K-long’ $K_{L}^{0}$ and short-lived ‘K-short’ $K_{S}^{0}$, which decay into three and two pions with a mean lifetimes of 15 m/$c$ and2.8 cm/$c$, respectively. By neglecting the small CP violation, the weak and strong kaonic eigenstates are related by:
@@
This prediction serves as the initial reference for testing the isospin symmetry with the results provided by the NA61/SHINE experiment [12] that measured $R_{K}=1.262\pm 0.017(stat)\pm 0.062(sys)$ in collisions of argon and scandium nuclei at the energy of 11.9 GeV per nucleon pair. This result is consistent with the compilation of the world data, even if each previous measurement, taken individually, was also consistent with $R_{K}=1$ within its large uncertainty. Yet, as discussed below, all data taken together clearly show a significant deviation from unity.
@@
Figure 1: The charged-to-neutral kaon ratio $R_{K}$ as a function of collision energy. The symbols show the experimental data. The black line shows the HRG predictions for $Q/B=0.4$. The blue dots indicate the HRG predictions for $Q/B$ values corresponding to the ones in experiments. For different nuclei, $Q/B$ corresponds to the electric charge over baryon number of the whole system. The grey rectangles show the UrQMD model results.
@@
The weighted average of all results is $1.143\pm 0.026$, where the second number stands for the error calculated following the Particle Data Group procedure [9]. The weighted average deviates from the expectation for the isospin symmetry broken by the known effects by $5.55$ errors. The $\chi_{\rm min}^{2}/\text{dof}=0.47$ may indicate an overestimation of the uncertainties. The estimated scaling factor for the uncertainty is 2.15, and the corresponding scaled error is 0.018. The weighted average deviates from the expectation for the isospin symmetry, broken by the known effects by 8.14 scaled errors.
@@
    An interesting experimental test is also possible by considering $\pi^{-}+C$ and $\pi^{+}+C$ interactions. While the ensemble of only one of them is not invariant under charge transformation, the ensemble having an equal number of $\pi^{-}+C$ and $\pi^{+}+C$ interactions is invariant. Thus, for the joint ensemble, the exact charge symmtery predicts $R_{K}=1$.
@@
    Kaons play a special role due to the simple isospin structure and easy-to-measure measure. This explains why the first results on a large isospin-symmetry breaking are reported for kaons. Yet, it is important to perform a similar study for other isospin multiplets in the future. For example, using the same methods, we have checked that the (anti-)proton to (anti)-neutron ratio is even less sensitive to the known isospin symmetry-breaking effects and, thus, is predicted to be almost exactly one in nucleus-nucleus collisions with $Q/B=0.5$ in a broad range of collision energies including the low energies. Considering collision-by-collision fluctuations of charged and neutral kaon numbers would also be interesting. A puzzling result was recently reported by ALICE at LHC [45] and discussed in Ref. [46].
@@
    The relationship between the production rates of charged and neutral kaons in hadronic collisions was used to estimate the flux of secondary neutral kaons produced in $p$+Be collisions at beam momenta of several 100 GeV/$c$ in neutral kaon and neutrino beams [47]. Recently, the charged-to-neutral kaon ratio was studied in p+p interactions at SPS energies, and no excess in comparison to a simple quark counting model emerged [48]. On the other hand, $\pi^{-}$+C interactions at SPS energies yield the ratio of charged-to-neutral kaons, which is significantly larger than unity and predictions of popular models [49].
@@ Methods
The UrQMD transport model [23, 24, 25] describes A+A collisions by explicitly propagating hadrons in phase space. During the propagation, rescattering among hadrons takes place. The particle production in this model happens either via resonance decay or via a string excitation and fragmentation following the LUND model [64]. The grey rectangles in Fig. 1 indicate the UrQMD model predictions. Here, we have considered central Au+Au collisions ($A=197,Z=79$, $Q/B\approx 0.4$). The predictions are shown within the $\sqrt{s_{NN}}$ range of 2.4 to 20 GeV. At each energy, $10^{4}$ events are used for the analysis.

### 01 Feb 2024 v2

@@ Abstract
The neutral kaons are eigenstates of strong interaction, oscillating ($K^{0}\leftrightarrow\bar{K}^{0}$) due to weak interaction. The weak eigenstates are the long-lived ‘K-long’ $K_{L}^{0}$ and short-lived ‘K-short’ $K_{S}^{0}$, which decay into three and two pions with mean lifetimes of 15 m/$c$ and 2.8 cm/$c$, respectively. By neglecting the small CP violation, the weak and strong kaonic eigenstates are related by:
@@
This prediction serves as the initial reference for testing the isospin symmetry with the results provided by the NA61/SHINE experiment [12] that measured $R_{K}=1.233\pm 0.014(stat)\pm 0.053(sys)$ in collisions of argon and scandium nuclei at the energy of 11.9 GeV per nucleon pair. This result is consistent with the compilation of the world data, even if each previous measurement, taken individually, was also consistent with $R_{K}=1$ within its large uncertainty. Yet, as discussed below, all data taken together clearly show a significant deviation from unity.
@@
Figure 1: The charged-to-neutral kaon ratio $R_{K}$ as a function of collision energy. The symbols show the experimental data. The black line shows the HRG predictions for $Q/B=0.4$. The blue dots indicate the HRG predictions for $Q/B$ values corresponding to the ones in experiments. For different nuclei, $Q/B$ corresponds to the electric charge over baryon number of the whole system. The grey squares show the UrQMD model results.
@@
The weighted average of all results is $1.144\pm 0.026$, where the second number stands for the error calculated following the Particle Data Group procedure [9]. The weighted average deviates from the expectation for the isospin symmetry broken by the known effects by $5.64$ errors. The $\chi_{\rm min}^{2}/\text{dof}=0.47$ may indicate an overestimation of the uncertainties. The estimated scaling factor for the uncertainty is 2.12, and the corresponding scaled error is 0.018. The weighted average deviates from the expectation for the isospin symmetry, broken by the known effects by 8.21 scaled errors.
@@
    An interesting experimental test is also possible by considering $\pi^{-}+C$ and $\pi^{+}+C$ interactions. While the ensemble of only one of them is not invariant under charge transformation, the ensemble having an equal number of $\pi^{-}+C$ and $\pi^{+}+C$ interactions is invariant. Thus, for the joint ensemble, the exact charge symmetry predicts $R_{K}=1$.
@@
    Kaons play a special role due to their simple isospin structure and easy measurement. This explains why the first results on a large isospin-symmetry breaking are reported for kaons. Yet, it is important to perform a similar study for other isospin multiplets in the future. For example, using the same methods, we have checked that the (anti-)proton to (anti)-neutron ratio is even less sensitive to the known isospin symmetry-breaking effects and, thus, is predicted to be almost exactly one in nucleus-nucleus collisions with $Q/B=0.5$ in a broad range of collision energies including the low energies. Considering collision-by-collision fluctuations of charged and neutral kaon numbers would also be interesting. A puzzling result was recently reported by ALICE at LHC [45] and discussed in Ref. [46].
@@
    The relationship between the production rates of charged and neutral kaons in hadronic collisions was used to estimate the flux of secondary neutral kaons produced in $p$+Be collisions at beam momenta of several 100 GeV/$c$ in neutral kaon and neutrino beams [47]. Recently, the charged-to-neutral kaon ratio was studied in p+p interactions at SPS energies, and no excess in comparison to a simple quark counting model emerged [48]. On the other hand, $\pi^{-}$+C interactions at SPS energies yield a ratio of charged-to-neutral kaons that is significantly larger than unity and predictions of popular models [49].
@@ Methods
The UrQMD transport model [23, 24, 25] describes A+A collisions by explicitly propagating hadrons in phase space. During the propagation, rescattering among hadrons takes place. The particle production in this model happens either via resonance decay or via a string excitation and fragmentation following the LUND model [64]. The grey squares in Fig. 1 indicate the UrQMD model predictions. Here, we have considered central Au+Au collisions ($A=197,Z=79$, $Q/B\approx 0.4$). The predictions are shown within the $\sqrt{s_{NN}}$ range of 2.4 to 20 GeV. At each energy, $10^{4}$ events are used for the analysis.